OpenStudy (anonymous):

8/121x + 9/11x^2 (those are fractions) sorry my brain just isn't working right now

OpenStudy (jdoe0001):

so... what do you think is the LCD anyhow?

OpenStudy (jdoe0001):

\(\bf \cfrac{8}{121x}+\cfrac{9}{11x^2}\implies \cfrac{8}{11\cdot 11x}+\cfrac{9}{11x\cdot x}=\cfrac{}{{\color{brown}{ lcd?}}}\) any ideas?

OpenStudy (anonymous):

(99x^2 + 8x)/121 You can simplify if you're supposed to

OpenStudy (michele_laino):

@M_e_g_a is your expression this? \[\frac{ 8 }{ 121 }x+\frac{ 9 }{ 11 }x ^{2}\]

OpenStudy (anonymous):

@Michele_Laino yes, the X's are on the denominators

OpenStudy (anonymous):

it doesn't matter where the x's are...

OpenStudy (anonymous):

@jdoe0001 The LCD would be 11x? or 11x^2...?

OpenStudy (jdoe0001):

the lcd, would be, one that comprises both 11x alone wouldn't

OpenStudy (michele_laino):

@M_e_g_a lcd is this: \[11^{2}*x ^{2}\]

OpenStudy (anonymous):

@jdoe0001 1+11^2+x?

OpenStudy (anonymous):

that one should be an eleven

OpenStudy (anonymous):

No i wrote that completely wrong, I meant to put 11+11x+x

OpenStudy (jdoe0001):

well... in this case one 11x doesn't by itself and \(11^2x\) doesn't by itself either but \(11^2x^2\) would keep in mind that, at the very least, one could use the product of all denominators as lcd

OpenStudy (michele_laino):

@M_e_g_a your result is: \[\frac{ 8x+11 }{ 121x ^{2} }\]

OpenStudy (anonymous):

but the 11x appears twice so you only have to use it once... that's what I thought

OpenStudy (michele_laino):

sorry I have made an error: \[\frac{ 8x+99 }{ 121x ^{2} }\]

OpenStudy (jdoe0001):

hmmm once you have the lcd... you'd add the fractions like you'd any other fraction divide the denominator by the lcd the quotient multiplies the numerator and adds to the next fraction result

OpenStudy (anonymous):

@Michele_Laino @jdoe0001 you guys were right I'm just a bit stupid xD thanks